1. A bank statement shows a balance of $6,825 at June 30. The bank reconciliation is prepared and includes outstanding checks of $2,850, deposits in transit of $1,250, and a bank service charge of $40. Among the paid checks returned by the bank was check number 900 in the amount of $450, which the company had erroneously recorded in the accounting records as $60. The "adjusted cash balance" at June 30 is:

Multiple Choice

  • $5,985.

  • $5,225.

  • $3,525.

  • $4,425.

2. The Cash account in the records of Hensley, Incorporated showed a balance of $4,300 at June 30. The bank statement, however, showed a balance of $6,700 at the same date. The only reconciling items consisted of a $750 deposit in transit, a bank service charge of $11, and a large number of outstanding checks.

What is the total amount of the outstanding checks at June 30?

Multiple Choice

  • $3,550.

  • $3,150.

  • $2,778

  • $3,161.
Answer #1

Solution 1:

adjusted cash balance = bank statement balance + deposit in transit - outstanding check

= 6825 + 1250 - 2850 = $5,225

Hence second option is correct.

Solution 2:

Adjusted balance = cash account balance - bank service charge = $4300 - $11 = $4,289

Outstanding checks = Bank balance + deposit in transit - Adjusted balance

= 6700 + 750 - 4289

= $3,161

Hence last option is correct.